# Towards Generalized Video Quality Assessment: A Weak-to-Strong Learning Paradigm

### v3

@@ 4.1.2 Ensembling Homogeneous Teachers
For video pair generation, given a pair $(x^{A},x^{B})$, each VQA model $f_{\text{weak},i}$ produces quality scores $\hat{y}^{A}_{i}$ and $\hat{y}^{B}_{i}$. We compute the mean scores $\overline{y}^{A}$ and $\overline{y}^{B}$ [^3] These weak models are trained on the same dataset and thus share the same score scale., and the score variances $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{y}^{A}-\overline{y}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$ with $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, labels are assigned according to the statistical significance thresholds in (Zhu et al., 2024): “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
4.1.3 Integrating Heterogeneous Teachers
Another complementary approach is to diversify the teacher models in order to enrich the supervision signals. In this work, we leverage synthetic distortion simulators as specialized VQA models, which do not require human annotations for training and can be easily scaled. Concretely, we introduce three categories of synthetic distortions to emulate typical real-world degradations: spatial distortions, temporal distortions, and streaming distortions. Spatial distortions include resolution downscaling, Gaussian blur, Gaussian noise, darkening, and brightening, simulating capture-related artifacts. Temporal distortions cover jitter and stuttering, which mimic playback issues often observed in practice. Streaming distortions involve H.264 and H.265 compression, capturing compression artifacts introduced by modern media delivery platforms. The detailed simulation procedures are provided in Appendix A.3.
@@ 4.1.3 Integrating Heterogeneous Teachers
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. Pairs with a severity difference |ij|>1|i-j|>1 are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with |ij|=1|i-j|=1 receive “better” or “worse”. The “similar” label is intentionally excluded, as ij=0i-j=0 implies identical videos.
Figure 6: The framework of our iterative weak-to-strong training strategy.
4.2 Iterative Weak-to-Strong Training Strategy
Within our W2S training framework, we have demonstrated that the student model can surpass its teacher models. This observation naturally motivates an iterative strategy: once a student model is trained, it can be promoted to act as a new teacher, thereby enabling another round of weak-to-strong training. Through such iterative cycles, the student progressively inherits knowledge from its predecessors while further enhancing its generalization capability. Therefore, we adopt this iterative paradigm to continually refine the student model.
From the data perspective, we expect the training samples in the next iteration to pose challenges beyond the capacity of the current teacher models, thereby further expanding the capability of the student. To this end, we introduce a difficult-sample selection strategy for both types of supervision signals in Section 4.1. Specifically, given a student model $f_{\text{w2s}}^{(i)}$ trained in the $i$-th iteration, the construction of difficult samples is straightforward for synthetic distortion pairs described in Section 4.1.2, since ground-truth labels can be directly derived from the distortion levels. We use $f_{\text{w2s}}^{(i)}$ to infer the relative quality of these pairs and select only those misclassified by the student as the training data for the $(i+1)$-th iteration.
While for the video pairs described in Section 4.1.2, no ground-truth labels are available. To address this, we adopt the group maximum differentiation (gMAD) competition framework (Ma et al., 2018) to select pairs that exhibit the largest disagreement between VQA models. Given the weak model set $\{f_{\text{weak}}^{j}\}_{j=1}^{N_{\text{weak}}}$ used to train $f_{\text{w2s}}^{(i)}$, we first partition the video pool $D_{\text{w2s}}^{(i+1)}$ into $\xi$ uniform quality levels based on the predictions of $f_{\text{weak}}^{j}$, within which videos are assumed to have similar perceptual quality. We then select pairs that are maximally differentiated by the trained student model $f_{\text{w2s}}^{(i)}$ while indistinguishable to the weak model $f_{\text{weak}}^{j}$ by
[…]
Moreover, we also reverse the roles of $f_{\text{weak}}^{j}$ and $f_{\text{w2s}}^{(i)}$ to capture cases where the student perceives similar quality but the weak model disagrees. This strategy systematically exploits the decision boundary mismatches between student and teacher models, generating informative and challenging samples that drive further improvements in next-round W2S training.
4.3 Training Strategy
We employ the standard cross-entropy loss as a baseline objective. However, weak annotations inevitably contain noise, and directly supervising the student with cross-entropy risks overfitting to erroneous labels. To mitigate this, we introduce an auxiliary confidence loss (Burns et al., 2023; Guo et al., 2024) that encourages the student to reinforce its own confident predictions, particularly when they diverge from weak labels. The overall objective is formulated as
$\mathcal{L}=(1-\lambda)\,\mathcal{L}_{\text{CE}}+\lambda\,\mathcal{L}_{\text{conf}},$ (2)
where $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss, $\mathcal{L}_{\text{conf}}$ the confidence loss, and $\lambda$ adaptively balances label reliability against model predictions. Details of the confidence loss are provided in Appendix C.2.2.
For training data, we construct a total of 700700k annotated video pairs using the procedure described in Section 4.1.2 and Section 4.1.3. These pairs are partitioned into three subsets of 500500k, 100100k, and 100100k, denoted as $D_{\text{w2s}}^{(1)}$, $D_{\text{w2s}}^{(2)}$, and $D_{\text{w2s}}^{(3)}$, corresponding to the three stages of iterative training. A detailed breakdown of the dataset, as well as the complete training setup, is provided in Appendix A.1 and Appendix C.2.1.
Table 1: Performance comparison with state-of-the-art methods. Best and second-best results are marked in bold and underline, respectively. “Overall” represents the weighted average results based on the number of videos in each dataset.
@@ 4.3 Training Strategy
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
State-of-the-art VQA Methods - teachers
MinimalisticVQA(VII) (Sun et al., 2024) 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) (Sun et al., 2024) 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA (Wu et al., 2022) 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER (Wu et al., 2023a) 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align (Wu et al., 2023b) 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
State-of-the-art VQA Methods - others
VQA2 (Jia et al., 2024) 0.878 0.872 0.794 0.821 0.881 0.880 0.785 0.830 0.811 0.823 0.847 0.854
VQAThinker (Cao et al., 2025) 0.883 0.880 0.798 0.834 0.881 0.884 0.808 0.847 0.860 0.863 0.855 0.866
Our Weak-to-Strong VQA Methods
(I): Ensembling homogeneous teachers 0.883 0.877 0.804 0.829 0.883 0.876 0.799 0.830 0.843 0.845 0.856 0.860
(II): (I) + Integrating heterogeneous teachers 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
(III): (II) + Confidence loss 0.885 0.881 0.803 0.831 0.890 0.891 0.797 0.833 0.849 0.856 0.857 0.865
(IV): (III) + Iterative stage W2S training 0.886 0.883 0.803 0.834 0.898 0.897 0.810 0.841 0.858 0.864 0.860 0.868
(V): (IV) + Iterative stage W2S training 0.893 0.889 0.807 0.837 0.902 0.901 0.818 0.846 0.852 0.858 0.865 0.872
@@
Methods SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
State-of-the-art VQA Methods - teachers
MinimalisticVQA(VII) (Sun et al., 2024) 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) (Sun et al., 2024) 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA (Wu et al., 2022) 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER (Wu et al., 2023a) 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align (Wu et al., 2023b) 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
State-of-the-art VQA Methods - others
VQA2 (Jia et al., 2024) 0.613 0.698 0.656 0.741 0.332 0.413 0.415 0.474 0.678 0.689 0.583 0.623
VQAThinker (Cao et al., 2025) 0.767 0.806 0.856 0.845 0.528 0.610 0.573 0.624 0.586 0.626 0.615 0.658
Our Weak-to-Strong VQA Methods
(I): Ensembling homogeneous teachers 0.688 0.756 0.769 0.808 0.456 0.497 0.455 0.502 0.649 0.682 0.602 0.643
(II): (I) + Integrating heterogeneous teachers 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
(III): (II) + Confidence loss 0.708 0.763 0.796 0.829 0.523 0.606 0.579 0.643 0.713 0.742 0.672 0.717
(IV): (III) + Iterative stage W2S training 0.711 0.770 0.807 0.831 0.606 0.678 0.657 0.737 0.759 0.782 0.722 0.765
(V): (IV) + Iterative stage W2S training 0.723 0.776 0.799 0.828 0.683 0.749 0.698 0.758 0.772 0.807 0.745 0.789
4.4 Experimental Results
We present the experimental results in Table 1, highlighting five progressively enhanced models of our method: models (I)–(III) incrementally add components in Stage 1, while model (IV) and model (V) introduce iterative training in Stage 2 and Stage 3, respectively. We analyze them from the following aspects:
Ensembling Homogeneous Teachers. Compared with single-teacher supervision, we find that ensembling multiple teachers yields stronger student models that outperform all individual teachers as well as their corresponding students. This result further highlights the weak-to-strong effect in VQA and shows that improving the quality of teacher supervision amplifies this effect, consistent with prior findings.
Integrating Heterogeneous Teachers. We incorporate synthetic distortion simulators as specialized VQA models to extend the capability of the teacher ensemble. With synthetic distortion pairs, the student model achieves consistent improvements across all benchmarks, yielding marginal gains on in-domain datasets and substantial enhancements on OOD benchmarks. These results demonstrate that incorporating diverse VQA models as teachers enables joint supervision that consistently fosters more generalizable quality assessment.
Confidence Loss. Incorporating $\mathcal{L}_{\text{conf}}$ yields clear gains on OOD datasets. This indicates that confidence loss mitigates the adverse impact of noisy weak labels and enables the student to reinforce its own reliable predictions.

### v4

@@ 3.2.1 Pseudo-labeling using VQA Models
For soft ranking, we further compute the score variance for the video pair $\sigma^{2}_{A}$ and $\sigma^{2}_{B}$. Assuming the quality difference $\Delta=\overline{j}^{A}-\overline{j}^{B}$ follows a Gaussian distribution $\mathcal{N}(\Delta;0,\sigma^{2}_{\Delta})$, where $\sigma_{\Delta}=\sqrt{\sigma^{2}_{A}+\sigma^{2}_{B}}$, we assign labels based on statistical significance thresholds adapted from [77]: pairs are labeled as “superior” if $\Delta>2\sigma_{\Delta}$, “better” if $\sigma_{\Delta}<\Delta\leq 2\sigma_{\Delta}$, “similar” if $\sigma_{\Delta}<\Delta\leq\sigma_{\Delta}$, “worse” if $2\sigma_{\Delta}<\Delta\leq-\sigma_{\Delta}$, and “inferior” if $\Delta\leq-2\sigma_{\Delta}$.
3.2.2 Quality Ranking via Distortion Simulations
Figure 3: Overall framework of the proposed LMM-PVQA model. Compared to the standard LMM, our visual feature extractor adopts a dual-branch design with an additional motion encoder and projector to capture temporal-related distortions. The model takes a video pair and a text prompt as inputs and generates a text output specifying which video has better quality.
We introduce synthetic distortions to simulate typical degradation that may occur in real-world scenarios, which are categorized into three types: spatial distortions, temporal distortions, and streaming distortions. Specifically, spatial distortions include resolution resizing, Gaussian blur, Gaussian noise, darkening, and brightening, which simulate capture-related artifacts. Temporal distortions consist of jitter and stuttering, mimicking playback issues commonly observed in practical settings. Streaming distortions involve H.264 and H.265 compression, reflecting compression artifacts introduced by streaming media platforms. We denote this subset of data as PLVD-DS. The detailed generation of synthetic distortions are provided in the supplement file.
@@ 3.2.2 Quality Ranking via Distortion Simulations
We leverage distortion severity levels (e.g., constant rate factor for compression) as pseudo-labels to infer relative quality. Given a primary video $x^{0}$ and a synthetic distortion simulator $\mathcal{S}$, we degrade $x^{0}$ across $N_{\mathcal{S}}$ severity levels to generate distorted videos $\{x_{\mathcal{S}}^{i}\}^{N_{\mathcal{S}}}_{i=1}$. Pairs $(x_{\mathcal{S}}^{i},x_{\mathcal{S}}^{j})$ are randomly sampled. For hard ranking, pairs are directly annotated as “better” if $i<j$ or “worse” otherwise. For soft ranking, pairs with a severity difference |ij|>1|i-j|>1 are labeled as “superior” or “inferior” depending on the relative order of $i$ and $j$, while pairs with |ij|=1|i-j|=1 receive “better” or “worse”. The “similar” label is intentionally excluded, as ij=0i-j=0 implies identical videos.
3.2.3 Label Refinement
The label quality of the PLVD-VQA dataset inherently depends on the performance of judges. To address this dependency, we introduce an iterative label refinement framework. Specifically, once a new VQA model is trained, we treat it as an improved judge $\mathcal{J}^{\prime}$ and reapply the annotation pipeline of PLVD-VQA to iteratively refine video pair labels.
In summary, the PLVD dataset comprises 700700k annotated video pairs, partitioned into three subsets of 500500k, 100100k, and 100100k pairs, denoted as PLVD-Part1, PLVD-Part2, PLVD-Part3. The latter two subsets are dedicated to iterative label refinement via the trained model. A detailed breakdown of PLVD, including pair types and the corresponding number of videos, is provided in Table 1.
4 Proposed Method
We aim to train a VQA model on an unlabeled video dataset to compute the perceptual quality score of a video. To achieve this goal, we reformulate the VQA regression task as a classification problem that distinguishes the relative quality between pairs of videos.
4.1 Model Structure
We introduce an LMM-based VQA framework for pairwise quality ranking (LMM-PVQA). As illustrated in Fig. 3, our model comprises three components: a visual feature extractor, a text tokenizer, and an LLM decoder.
$\bm{r}$ $=\mathcal{L}(\bm{f}^{v}_{A},\bm{f}^{v}_{B},\bm{f}^{p}),$ (2)
where $\bm{r}$ is expected to belong to {“better”, “worse”} for hard ranking and {“superior”, “better”, “similar”, “worse”, “inferior”} for soft ranking.
Table 2: Performance comparison of our models against competitive opinion-unaware and supervised methods in the zero-shot setting. The best results are highlight in boldface, the second best is underlined. NA in the table indicates unavailable results. “Overall” represents the weighted average results based on the number of videos in each dataset. PLVD-P1/P2/P3 represent PLVD-Part1/ Part2/ Part3 respectively, A ✔ in the ”Label” column indicates that human-labeled data is used for training
@@ 4.1 Model Structure
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.442 0.332 0.489 0.459 0.541 0.553 0.596 0.628 0.278 0.290 0.457 0.395
IL-NIQE [72] 0.483 0.362 0.418 0.397 0.512 0.530 0.484 0.532 0.291 0.323 0.454 0.390
VIIDEO [39] 0.080 0.080 0.009 0.019 0.299 0.300 0.033 0.215 0.058 0.154 0.077 0.095
STEM [20] 0.206 0.243 0.434 0.381 0.619 0.627 0.594 0.629 0.284 0.318 0.325 0.336
TPQI [27] NA NA NA NA 0.556 0.549 0.636 0.645 0.111 0.218 0.411 0.449
BUONA-VISTA [63] NA NA NA NA 0.760 0.760 0.784 0.794 0.525 0.556 0.680 0.693
Supervised VQA Methods
MinimalisticVQA(VII) [52] LSVQ [70] 0.861 0.859 0.740 0.784 0.843 0.841 0.757 0.813 0.775 0.779 0.817 0.830
MinimalisticVQA(IX) [52] LSVQ [70] 0.885 0.882 0.792 0.828 0.862 0.859 0.775 0.821 0.826 0.821 0.849 0.859
FAST-VQA [61] LSVQ [70] 0.880 0.880 0.781 0.813 0.859 0.854 0.826 0.845 0.730 0.747 0.838 0.849
DOVER [64] LSVQ [70] 0.878 0.866 0.782 0.813 0.874 0.869 0.817 0.840 0.771 0.781 0.842 0.845
Q-Align [65] fused [17, 11, 28, 70, 40] 0.886 0.884 0.761 0.822 0.876 0.878 0.783 0.819 0.834 0.846 0.844 0.861
Our Self-supervised VQA Methods: LMM-PVQA
Hard Ranking PLVD-P1 (500k) 0.883 0.866 0.799 0.817 0.886 0.869 0.791 0.820 0.839 0.844 0.854 0.850
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.888 0.884 0.806 0.835 0.894 0.893 0.801 0.836 0.854 0.861 0.861 0.868
@@
Methods Training data Label SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
Traning-free / Opinion-unaware VQA Methods
NIQE [38] None 0.240 0.247 0.473 0.496 0.354 0.413 0.048 0.002 0.163 0.114 0.183 0.154
IL-NIQE [72] 0.200 0.168 0.340 0.303 -0.081 -0.040 0.079 -0.008 0.056 0.006 0.083 0.036
VIIDEO [39] 0.077 -0.199 0.157 -0.257 0.276 0.244 0.114 0.078 0.082 0.019 0.110 0.010
STEM [20] 0.103 0.111 0.498 0.492 0.288 0.317 0.184 0.097 0.123 0.104 0.172 0.147
TPQI [27] NA NA NA NA NA NA NA NA NA NA NA NA
BUONA-VISTA [63] NA NA NA NA NA NA NA NA NA NA NA NA
Supervised VQA Methods
MinimalisticVQA(VII) [52] LSVQ [70] 0.596 0.682 0.681 0.733 0.061 0.130 0.275 0.338 0.604 0.659 0.490 0.551
MinimalisticVQA(IX) [52] LSVQ [70] 0.686 0.746 0.797 0.816 0.301 0.388 0.459 0.502 0.615 0.661 0.574 0.622
FAST-VQA [61] LSVQ [70] 0.631 0.677 0.725 0.747 0.326 0.415 0.327 0.363 0.518 0.526 0.486 0.512
DOVER [64] LSVQ [70] [40] 0.647 0.728 0.694 0.747 0.360 0.465 0.368 0.418 0.559 0.593 0.519 0.569
Q-Align [65] fused [17, 11, 28, 70, 40] 0.611 0.681 0.756 0.798 0.329 0.342 0.414 0.497 0.613 0.655 0.555 0.606
Our Self-supervised VQA Methods: LMM-PVQA
Hard Ranking PLVD-P1 (500k) 0.705 0.742 0.794 0.801 0.496 0.550 0.492 0.542 0.640 0.670 0.613 0.648
Soft Ranking (Stage 1) PLVD-P1 (500k) 0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
Soft Ranking (Stage 2) PLVD-P1/ P2 (600k) 0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
Soft Ranking (Stage 3) PLVD-P1/ P2/ P3 (700k) 0.703 0.761 0.792 0.823 0.559 0.644 0.670 0.728 0.754 0.770 0.716 0.752
4.1.1 Training and Inference Pipelines
Training. LMM-PVQA is optimized via standard supervised fine-tuning (SFT). To enhance training efficiency, the LLM parameters are kept frozen to preserve their intrinsic reasoning capabilities. Similarly, the pretrained SlowFast motion encoder remains frozen due to its proven effectiveness in capturing temporal representations [52]. Trainable parameters are confined to the image encoder $\mathcal{F_{I}}$ and feature projection layers $\mathcal{P}_{I}$ and $\mathcal{P}_{M}$, thus enabling domain-specific visual feature adaptation.
Furthermore, we propose an iterative self-improvement training strategy. The training process consists of three stages. First, the base model is trained on the PLVD-Part1 dataset. Next, the trained model acts as an enhanced judge, annotating video pairs in the PLVD-Part2 dataset alongside existing judges. The combined PLVD-Part1 and PLVD-Part2 datasets are then used to train a second-stage model. This process is repeated for a third-stage model. As demonstrated in Section 5.2, this iterative strategy progressively enhances out-of-distribution performance.
Inference. Given that the model primarily predicts relative quality rankings between video pairs, a practical conversion to absolute quality scores is required. To achieve this, we adopt an adaptive soft comparison method proposed in  [77], which first computes a soft probability matrix across ranking categories by comparing the test video against anchor videos, followed by maximum a posteriori (MAP) [54] estimation under Thurstone’s Case V model [53] to obtain calibrated quality scores.
For anchor selection, the range of quality scores of PLVD-VQA is partitioned into five quality intervals. Within each interval, the video exhibiting the lowest inter-model score variance is selected as the anchor, ensuring stable reference points for score derivation. We list the detailed calculation procedure in the supplement file.
5 Experiments
5.1 Experimental Setups
Implementation Details For each stage of training, we initialize the image encoder, image projector, and LLM with LLaVA-ov-chat (7B) weights [24] and the motion encoder with SlowFast pre-trained weights. The model is trained for one epoch on four NVIDIA A800 GPUs with a total batch size of $4$ and a learning rate of 11e-55. Inference is performed on two NVIDIA RTX 3090 GPUs.
Benchmark Datasets The proposed methods is evaluated on ten VQA benchmarks, categorized into in-domain and out-of-distribution (OOD) groups to systematically assess their generalization. The in-domain evaluation benchmarks include LSVQ Test [70], LSVQ 1080p [70], KoNViD-1k [16], LIVE-VQC [50], and YouTube-UGC [60], all of which contain UGC videos. The OOD datasets consist of LIVE-YT-Gaming [71], CGVDS [47], LIVE-YT-HFR [33], Waterloo-IVC-4K [26], and KVQ [32]. Specifically, LIVE-YT-Gaming and CGVDS contain gaming videos, which exhibit significant differences in content compared to natural UGC videos. LIVE-YT-HFR includes videos with varying frame rates, focusing on temporal-related distortions. Waterloo-IVC-4K and KVQ primarily evaluate compression artifacts, where Waterloo-IVC-4K contains compressed high-resolution videos at different spatial resolutions, while KVQ comprises compressed UGC videos processed with various video enhancement techniques. We introduce details of these datasets in the supplement file.
Evaluation Criteria We adopt two widely used criteria to evaluate the performance of VQA models: Spearman Rank Correlation (SRCC) and Pearson Linear Correlation (PLCC), which indicate the prediction monotonicity and prediction linearity, respectively.
Competing Methods We compare our model with six opinion-unaware BVQA methods, including NIQE [38], IL-NIQE [72], VIIDEO [39], STEM [20], TPQI [27], and BUONA-VISTA [63], as well as five supervised VQA methods: Minimalistic-VQA (VII) [52], Minimalistic-VQA (IX) [52], FAST-VQA [61], DOVER [64], and Q-Align [65]. These five supervised VQA models serve as strong baselines, as they are used to pseudo-label video quality and can be regarded as teacher models. The supervised models are trained on the LSVQ training set.
Table 3: Supervised performance comparison of our model against competitive methods. The experimental results on KoNViD-1k, YouTube-UGC, and LIVE-VQC are obtained using 10-fold cross-validation. “Base” denotes training LMM-PVQA without self-supervised pretraining
Testing Set $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC
TLVQM [22] 0.772 / 0.774 0.589 / 0.616 0.732 / 0.724 0.670 / 0.691 0.685 / 0.692
VIDEAL [55] 0.794 / 0.783 0.545 / 0.554 0.751 / 0.741 0.630 /0.640 0.687 / 0.709
VSFA [25] 0.801 / 0.796 0.675 / 0.704 0.794 / 0.799 0.718 / 0.771 0.787 / 0.789
PatchVQ [70] 0.827 / 0.828 0.711 / 0.739 0.791 / 0.786 0.827 / 0.837 NA / NA
SimpleVQA [51] 0.867 / 0.861 0.764 / 0.803 0.856 / 0.860 0.845 / 0.859 0.847 / 0.856
FAST-VQA [61] 0.880 / 0.880 0.781 / 0.813 0.891 / 0.892 0.849 / 0.865 0.855 / 0.852
DOVER [64] 0.878 / 0.866 0.782 / 0.813 0.908 / 0.910 0.860 / 0.875 0.841 / 0.851
MinimalisticVQA [52] 0.881 / 0.879 0.781 / 0.820 0.889 / 0.890 0.842 / 0.854 0.890 / 0.891
Soft Ranking (Base) 0.880 / 0.864 0.790 / 0.814 0.896 / 0.877 0.840 / 0.853 0.858 / 0.848
Soft Ranking (Stage 1) 0.907 / 0.904 0.832 / 0.857 0.911 / 0.908 0.884 / 0.894 0.910 / 0.911
Table 4: Experimental results of ablation study. “Spatial” and “Motion” denote the spatial branch and the temporal branch of the visual feature extractor. In Stage 1, if PLVD-DS (Part1) is not specified, the model is trained solely on PLVD-VQA (Part1). In Stage 2, “w/o label refinement” and “w/ label refinement” indicate whether the trained model is used to label the data in PLVD-Part2
Stage 1 Stage 2 $_{\text{test}} LSVQ$$_{\text{1080p}}$ KoNViD-1k LIVE-VQC YouTube-UGC Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.872 0.862 0.787 0.814 0.868 0.864 0.783 0.819 0.834 0.839 0.843 0.846
0.883 0.872 0.804 0.827 0.883 0.876 0.799 0.830 0.842 0.843 0.855 0.857
0.886 0.880 0.803 0.830 0.891 0.888 0.797 0.832 0.845 0.849 0.858 0.863
0.887 0.879 0.801 0.829 0.890 0.884 0.791 0.826 0.853 0.855 0.858 0.862
0.887 0.880 0.802 0.830 0.893 0.892 0.798 0.833 0.856 0.859 0.859 0.864
Stage 1 Stage 2 LIVE-YT-Gaming CGVDS LIVE-YT-HFR Waterloo-IVC-4K KVQ Overall
Spatial Motion PLVD-DS w/o label refinement w/ label refinement SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
0.678 0.735 0.754 0.800 0.334 0.413 0.404 0.459 0.616 0.656 0.561 0.610
0.688 0.736 0.769 0.808 0.446 0.477 0.425 0.470 0.619 0.662 0.579 0.622
0.697 0.752 0.799 0.829 0.481 0.525 0.552 0.614 0.690 0.725 0.650 0.693
0.688 0.741 0.794 0.822 0.446 0.508 0.541 0.608 0.705 0.735 0.651 0.694
0.717 0.763 0.810 0.834 0.515 0.611 0.633 0.696 0.749 0.764 0.704 0.741
5.2 Performance Analysis
The performance of all compared methods and our proposed models is summarized in Table 2, with detailed analysis conducted from the following perspectives:
Hard Ranking v.s. Soft Ranking. The model trained with soft ranking marginally outperforms its hard-ranking counterpart on in-domain VQA benchmarks and CGVDS. Moreover, it achieves significantly better performance on Waterloo-IVC-4K and KVQ, while exhibiting slightly inferior results on LIVE-YT-Gaming and LIVE-YT-HFR. Overall, the soft ranking strategy incorporates “quality distance” information into the training process, enabling the model to easily learn quality comparison capabilities and achieve better performance.